Please help me to solve a problem by using a proportion:
66) Oil and Gas: An outboard motor calls for a fuel mixture that has a gasoline-to-oil ratio of 50 to 1. How many pints of oil should be added to 6 gallons of gasoline?

Please help me to solve each problem:
76) Magazine Sales. A company specializing in magazine sales over the phone found that in 2500 phone calls, 360 resulted in sales and were made by male callers, and 480 resulted in sales and were made by female callers. If the company gets twice as many sales per call with a woman's voice that with a man's voice, then how many of the 2500 calls were made by females?
Thank you very much for your time and help!

Answer by stanbon(75887)

66) Oil and Gas: An outboard motor calls for a fuel mixture that has a gasoline-to-oil ratio of 50 to 1. How many pints of oil should be added to 6 gallons of gasoline?
1/50 = x/6
x= 6/50
x= 3/25
x = 0.12 pints of oil for 6 gallons of gasoline
-------------------------
Please help me to solve each problem:
76) Magazine Sales. A company specializing in magazine sales over the phone found that in 2500 phone calls, 360 resulted in sales and were made by male callers, and 480 resulted in sales and were made by female callers. If the company gets twice as many sales per call with a woman's voice than with a man's voice, then how many of the 2500 calls were made by females?
-----
Let "x" be the number of men's calls; Rate of men's sales = 360/x
2500-x is the number of women's calls: Rate of women's sales = 480/(2500-x)
------
EQUATION:
Rate of women's sales = 2*rate of men's sales
480/(2500-x) = 2(360/x)
2/(2500-x) = 3/x
2x = 7500-3x
5x = 7500
x = 1500 (number of men's calls)
2500-x=1000(number of women's calls)
=============
Cheers,
Stan H.